# SHAM-OT: Rapid Subhalo Abundance Matching with Optimal Transport

### v1

@@ 1 Introduction
To create the SHAM matching function, the galaxy and halo distributions are represented by samples; a large number of samples is required for the matching to be accurate. This procedure is typically repeated at multiple redshifts. The generation and sorting of these samples may become computationally expensive for large catalogues, in terms of both time and memory. While this is manageable for single simulations, it may become a bottleneck for Bayesian inference that marginalizes over uncertainties in the stellar mass or luminosity functions, as the matching has to be re-computed for any new combination of parameters. Another important element of SHAM is the scatter in the galaxy-halo relation; its inclusion should preserve the halo mass and galaxy luminosity functions. For certain models, such as log-normal stellar mass scatter (Behroozi et al. (2010), hereafter B10), the “original” stellar mass function is computed by deconvolving its “noisy” counterpart with the scatter model.
In this work, we present SHAM-OT: a novel framework that reformulates the SHAM problem as optimal transport (OT). Generally, OT finds a 2D joint probability distribution (or a transport plan) such that its marginals are equal to two given 1D distributions, while minimising the cost of transporting one into the other, according to a given cost function (see Villani, 2009, for review). We show that the SHAM procedure is exactly the same as the solution to the optimal transport problem of empirical distributions (represented by samples), under certain assumptions on the transport cost function. Framing SHAM as OT allows us to operate directly in the large sample limit and use continuous PDF representations of the distributions, instead of samples. Thus, standard OT solvers can be used to find the matching from discretised halo and galaxy mass functions, without the need of creating catalogues. In this work we show the benefits of using the SHAM-OT framework: (i) computational efficiency, (ii) natural inclusion of scatter in the matching relation via OT plan regularization, and (iii) straightforward extension to matching more than two distributions using multi-marginal OT.
@@ 2 Optimal Transport
Figure 1: Analytical halo mass and luminosity function and the corresponding galaxy-halo connection derived using SHAM-OT and standard SHAM at redshift $z\!=\!0.5$. The standard SHAM approach can only be applied up to a halo mass of approximately $10^{14}M_{\odot}$ due to limited number of massive objects in the sampled catalogue, which here contains $\sim 10^{8}$ objects. In contrast, SHAM-OT operates directly on marginal functions, removing the need for sampling and sorting.
@@ 3 The Subhalo Abundance Matching Problem
Abundance matching techniques find the relationship between galaxies and dark matter haloes. In their simplest form, these methods assume a monotonic relationship between the halo mass and a given galaxy property, such as stellar mass or absolute magnitude. The galaxy-halo connection is then obtained by ranking both galaxies and haloes by these properties, and matching the most massive (or brightest) galaxies with the most massive haloes. Subhalo abundance matching (SHAM) extends this method to account for subhaloes and produces a relation that agrees better with observations (Kravtsov et al., 2004; Conroy et al., 2006; Vale & Ostriker, 2004). Additionally, more realistic models incorporate scatter in the galaxy-halo relation (Behroozi et al., 2010; Trujillo-Gomez et al., 2011; Reddick et al., 2013; Zentner et al., 2014). For example, B10 demonstrates how log-normal scatter of stellar mass at a fixed halo mass can be added, while preserving the marginal distribution of the stellar mass function. Furthermore, differentiating between passive red galaxies and star-forming blue galaxies enhances modelling accuracy, as these populations exhibit distinct preferences for occupation in haloes versus subhaloes (Dragomir et al., 2018; Rodríguez-Puebla et al., 2015; Yamamoto et al., 2015; Berner et al., 2024).
@@ 4 SHAM as Optimal Transport
Intuitively, for the sorted matched samples, the cost cannot be lowered by exchanging any two matches. Theorem 4.3 in Bobkov & Ledoux (2019) establishes the approximation bounds for transport cost between the empirical distribution and the true distribution. By the Fundamental Theorem of Statistics (Glivenko–Cantelli), the empirical distribution converges to the true distribution $\pi^{*}_{N}\to\pi^{*}$ when $N\to\infty$ almost surely. Therefore, in the limit of $N\to\infty$, the solution to the SHAM procedure is equal to the solution of the continuous optimal transport in Eqn. 2 with any metric cost function. The regularized OT, however, will depend on the choice of the cost.
Figure 3: Comparison of scatter models: SHAM-OT with entropic regularization and standard SHAM. Left: Mean magnitude as a function of halo mass for different regularization $\epsilon$ for SHAM-OT and scatter parameters $\sigma_{\mathrm{mag}}$ for the standard SHAM. Right: Conditional transport plan for $\epsilon\!=\!3\!\times\!10^{-6}$ and $\sigma_{\mathrm{mag}}\!=\!0.6$. The cost is defined using $\mu\!\propto\!\log_{10}M$ and $\nu\!\propto\!-\!m$ (see Sec. 5).
Figure 4: Impact of different cost functions on the conditional transport plan, for quadratic (left) and absolute difference (right) costs, with $\epsilon\!=\!0.01$. Quadratic cost results in increased scatter for lower masses. Absolute difference cost reduces the scatter for low masses. This demonstrates how regularization can serve as a flexible tool to model different types of noise behaviour.
@@ 5 Matching halo mass and galaxy luminosity with Optimal Transport
Popular choices for the galaxy mass proxy include the stellar mass and the absolute magnitude (see Wechsler & Tinker (2018) for an overview). In this work, we use the absolute magnitude in the B-band in the parametrization from Fischbacher et al. (2024a). The number of galaxies $n_{\mathrm{g}}$ per absolute magnitude $m$ is given by
@@
where $M^{*}_{1},M_{2}^{*},\phi^{*}_{1},\phi^{*}_{2}$ and $z_{0}$ are free parameters. The free parameters are fixed to the first index of the blue population of the publicly available model by Fischbacher et al. (2024a).
@@
The transport of distributions requires them to have the same total mass. We ensure this by defining a lower limit on the host mass: we choose $M^{\mathrm{min}}\!=\!10^{10}\!M_{\odot}$, and then compute the necessary faintest absolute magnitude such that the number of haloes and galaxies per unit volume is equal. For numerical reasons, we also define an upper limit on both the halo mass ($M^{\mathrm{max}}{=}10^{19}M_{\odot}$) and the brightest absolute magnitude ($m^{\mathrm{min}}$=-27). These numbers are chosen such that the number of haloes or galaxies above those limits is so small that it does not impact the matching results from SHAM. We create a linearly-spaced vector of negative magnitudes $\tilde{\bm{\nu}}\!\in\!\mathbb{R}^{n}\!:\!\tilde{\nu}_{i}\!=\!-m_{i}$ and logarithmically-spaced halo mass vector $\tilde{\bm{\mu}}\!\in\!\mathbb{R}^{n}\!:\!\tilde{\mu}_{i}\!=\!\log_{10}M_{i}$, both with $n\!=\!500$ within the limits described above, resulting in sampling spacing of $\Delta\tilde{\nu}\!=\!0.036$ and $\Delta\tilde{\mu}\!=\!0.018$. We generate the discretized halo mass function $\mathbf{p}\!\in\!\mathbb{R}_{+}^{n}$ using equations 9 and 11, and the galaxy absolute magnitude function $\mathbf{q}\in\mathbb{R}_{+}^{n}$ using Eqn. 12. We show $\mathbf{p}$ and $\mathbf{q}$ in side panels in Fig. 1. To calculate the cost, we define $\bm{\nu}\!=\!(\tilde{\bm{\nu}}\!-\!\tilde{\nu}_{\mathrm{ref}})/\Delta\tilde{\nu}$ and $\bm{\mu}\!=\!(\tilde{\bm{\mu}}\!-\!\tilde{\mu}_{\mathrm{ref}})/\Delta\tilde{\mu}$, with reference values $\tilde{\nu}_{\mathrm{ref}}\!=\!9.98$ and $\mu_{\mathrm{ref}}\!=\!9.09$. We define the cost matrix $\mathbf{C}$ as $C_{ij}\!=\!(\nu_{i}\!-\!\mu_{j})^{2}/\mu_{j}^{2}$. We obtain the optimal transport plan by solving the problem in Eqn. 4 using linear programming. In the standard SHAM approach, we sample catalogues in a redshift interval $z\!\in\![0.495,0.505]$, which is typical for practical abundance matching, resulting in $\sim\!10^{8}$ objects. The galaxy-halo connection can only be computed up to a mass limit of about $10^{14}M_{\odot}$, as the rapid decline of the mass functions prevents any higher-mass haloes in the simulated catalogue.
@@
The resulting optimal transport plan $\mathbf{Q}$ is presented in Fig. 1 and compared with the standard SHAM approach: they align almost perfectly. The residual differences stem from the fact that the standard SHAM is affected by noise from finite number of pairs, especially at the high-mass end. SHAM-OT overcomes this limitation by acting directly on the discretized functions, which can be extended to arbitrarily large halo masses. Furthermore, we find a computational speed improvement of $O(100)$ for SHAM-OT compared to standard SHAM. SHAM-OT’s computational efficiency allows us to easily study the impact of changes in the luminosity or halo mass function. In Fig. 2, we show how different values of the cosmological and luminosity function parameters impact the galaxy-halo connection.
@@ 6 Entropic Regularization as a Scatter Model
In the OT framework, the marginals of the transport plan are guaranteed to match the input distributions. The effect of scatter can be modelled naturally by the use of regularization of the transport plan: increasing the regularization parameter $\epsilon$ will result in a smoother plan. Various noise model behaviours, including the dependence on halo mass or galaxy luminosity, can be designed through different choices of the cost function, as long as it satisfies the metric requirements, as discussed in Sec. 4.
We demonstrate the capacity of the entropic regularization to model various scattering properties, including ones that mimic the log-normal approach. We use the same analytical prescription for the halo mass and luminosity function as described in Sec. 5. We compute the standard SHAM solution with Gaussian scatter in magnitude for three noise levels, using the method from B10 for three values of scatter $\sigma_{\mathrm{mag}}$. In SHAM-OT, we choose the cost matrix $C_{ij}\!=\!(\nu_{i}\!-\!\mu_{j})^{2}/\mu_{j}^{2}$. Three values of $\epsilon$ and the form of the cost function were chosen so that the resulting scatter closely resembles the Gaussian scatter of the standard approach. As before, we obtain the OT solution with Sinkhorn (Alg. 1). Fig. 4 shows the comparison between the two approaches for noise levels. SHAM-OT matches the log-normal model well, both in terms of the mean of the distribution (left panel) and its shape (right panel). It reproduces the trend that higher scatter shifts the mean transport plan from high halo masses to fainter galaxies and finds similar probability distributions.
Beyond emulating the log-normal scatter, regularized OT offers a flexible framework to model different scatter behaviours. Fig. 4 illustrates this flexibility by comparing the conditional transport plan for two different choices of cost functions. Further exploration of cost functions, along with the development of their physically motivation, is left for future work.
@@ 7 SHAM-OT with haloes from simulations
In practice, the halo mass function is often not analytically given, but can instead be computed from empirical halo distributions based on simulations. We use the approximate code PINOCCHIO (Monaco et al., 2002a; Taffoni et al., 2002; Monaco et al., 2002b, 2013; Munari et al., 2017; Rizzo et al., 2017) to simulate the past light cone of dark matter haloes. PINOCCHIO uses Lagrangian perturbation theory as an approximation which makes it faster than full N-body simulations. We assign subhaloes to the halo catalogue based on the merger tree using the method described in Berner et al. (2022) and label the haloes and subhaloes as hosts for red or blue galaxies based on the model from Berner et al. (2024). This model is based on two key parameters: the halo mass limit $M_{\text{limit}}\!=\!8\!\times\!10^{12}\,h^{-1}M_{\odot}$ and the satellite quenching time $t_{\text{quench}}\!=\!2.0$ Gyr. Central haloes with masses above $M_{\text{limit}}$ host red galaxies, while lower-mass haloes host blue galaxies. Galaxies in subhaloes remain blue until they have spent more than $t_{\text{quench}}$ within their host, after which they transition to red. As a result, we obtain catalogues of haloes and subhaloes with their corresponding mass, position in 3D space, and population label (i.e. red or blue). Note that the methods above can easily be replaced by any other approach that provides a halo catalogue containing the same parameters. We split the host catalogue into redshift bins with width $\Delta z\!=\!0.01$. For each bin, we obtain a host mass function as a histogram, including the upper (heaviest host in the redshift bin) and lower (mass resolution of the simulation) mass limits. We then utilize the luminosity function as defined in Fischbacher et al. (2024a) with the parameters for the red and blue population from the first index of the posterior distribution. Similar to the procedure described in Sec. 5, we determine the faintest absolute magnitude required such that the number of hosts equals the expected number of galaxies in this redshift bin. Using the constructed distributions for the host mass and galaxy absolute magnitudes, we compute the OT solution for each redshift bin, for both red and blue galaxies. The resulting galaxy-halo relation at different redshifts, for both red and blue galaxies, is shown in Fig. 5.

### v2

@@ 1 Introduction
To create the SHAM matching function, the galaxy and halo distributions are represented by samples; a large number of samples is required for the matching to be accurate. This procedure is typically repeated at multiple redshifts. The generation and sorting of these samples may become computationally expensive for large catalogues, in terms of both time and memory. While this is manageable for single simulations, it may become a bottleneck for Bayesian inference that marginalizes over uncertainties in the stellar mass or luminosity functions, as the matching has to be re-computed for any new combination of parameters. Another important element of SHAM is the scatter in the galaxy-halo relation; its inclusion should preserve the halo mass and galaxy luminosity functions. For certain models, such as log-normal stellar mass scatter (Behroozi et al., 2010, hereafter B10), the “original” stellar mass function is computed by deconvolving its “noisy” counterpart with the scatter model.
In this work, we present SHAM-OT: a novel framework that reformulates the SHAM problem as optimal transport (OT). Generally, OT finds a 2D joint probability distribution (or a transport plan) such that its marginals are equal to two given 1D distributions, while minimising the cost of transporting one into the other, according to a given cost function (see Villani, 2009, for review). We show that the SHAM procedure is exactly the same as the solution to the optimal transport problem of empirical distributions (represented by samples), under certain assumptions on the transport cost function. Framing SHAM as OT allows us to operate directly in the large sample limit and use continuous PDF representations of the distributions, instead of samples. Standard OT solvers can be used to find the matching from discretised halo and galaxy mass functions (represented as binned histograms, typically with $<\!100$ bins), avoiding the need to process large simulation catalogues (typically $>\!10^{6}$). In this work we show the benefits of using the SHAM-OT framework: (i) computational efficiency, (ii) natural inclusion of scatter in the matching relation via OT plan regularization, and (iii) straightforward extension to matching more than two distributions using multi-marginal OT.
@@ 2 Optimal Transport
Figure 1: Analytical halo mass and luminosity function and the corresponding galaxy-halo connection derived using SHAM-OT and standard SHAM at redshift $z\!=\!0.5$. The standard SHAM approach can only be applied up to a halo mass of approximately $10^{14}M_{\odot}$ due to limited number of massive objects in the sampled catalogue, which here contains $\sim 10^{8}$ objects. In contrast, SHAM-OT operates directly on binned mass and luminosity functions, removing the need for sampling and sorting.
@@ 3 The Subhalo Abundance Matching Problem
Abundance matching techniques find the relationship between galaxies and dark matter haloes. In their simplest form, these methods assume a monotonic relationship between the halo mass and a given galaxy property, such as stellar mass or absolute magnitude. The galaxy-halo connection is then obtained by ranking both galaxies and haloes by these properties, and matching the most massive (or brightest) galaxies with the most massive haloes. Subhalo abundance matching (SHAM) extends this method to account for subhaloes and produces a relation that agrees better with observations (Kravtsov et al., 2004; Conroy et al., 2006; Vale & Ostriker, 2004). Additionally, more realistic models incorporate scatter in the galaxy-halo relation (Behroozi et al., 2010; Trujillo-Gomez et al., 2011; Reddick et al., 2013; Zentner et al., 2014). For example, B10 demonstrates how log-normal scatter of stellar mass at a fixed halo mass can be added, while preserving the marginal distribution of the stellar mass function. Furthermore, differentiating between passive red galaxies and star-forming blue galaxies enhances modelling accuracy, as these populations exhibit distinct preferences for occupation in haloes versus subhaloes (Dragomir et al., 2018; Rodríguez-Puebla et al., 2015; Masaki et al., 2013; Yamamoto et al., 2015; Berner et al., 2024).
@@ 4 SHAM as Optimal Transport
Intuitively, for the sorted matched samples, the cost cannot be lowered by exchanging any two matches. Theorem 4.3 in Bobkov & Ledoux (2019) establishes the approximation bounds for transport cost between the empirical distribution and the true distribution. By the Fundamental Theorem of Statistics (Glivenko–Cantelli), the empirical distribution converges to the true distribution $\pi^{*}_{N}\to\pi^{*}$ when $N\to\infty$ almost surely. Therefore, in the limit of $N\to\infty$, the solution to the SHAM procedure is equal to the solution of the continuous optimal transport in Eqn. 2 with any metric cost function.
Figure 3: Comparison of SHAM-OT with entropic regularization and standard SHAM. Left: Mean magnitude vs. halo mass for different $\epsilon$ and $\sigma_{\mathrm{mag}}$. Right: Transport plan $\mathbf{Q}$ normalized by halo marginal $\mathbf{p}$, shown for the middle case on the left. SHAM-OT (blue) agrees well with standard SHAM (black).
Figure 4: Impact of cost function choices on the transport plan. The regularization parameter $\epsilon$ is tuned to reproduce the same scatter at $M=10^{12}M_{\odot}$ as in Figure 4. Left: Using exponent $\alpha=-4$ concentrates the plan at low halo masses. Right: Setting $\mu_{\mathrm{sc}}=0$ increases scatter at low masses.
@@ 5 Matching halo mass and galaxy luminosity with Optimal Transport
Popular choices for the galaxy mass proxy include the stellar mass and the absolute magnitude (see Wechsler & Tinker (2018) for an overview). In this work, we use the absolute magnitude in the B-band in the parametrization from Fischbacher et al. (2025) (hereafter F25). The number of galaxies $n_{\mathrm{g}}$ per absolute magnitude $m$ is given by
@@
where $M^{*}_{1},M_{2}^{*},\phi^{*}_{1},\phi^{*}_{2}$ and $z_{0}$ are free parameters. The free parameters are fixed to the first index of the blue population of the publicly available model by F25.
@@
The transport of distributions requires them to have the same total mass. We ensure this by defining a lower limit on the host mass: we choose $M^{\mathrm{min}}\!=\!10^{10}\!M_{\odot}$, and then compute the necessary faintest absolute magnitude such that the number of haloes and galaxies per unit volume is equal. For numerical reasons, we also define an upper limit on both the halo mass ($M^{\mathrm{max}}{=}10^{19}M_{\odot}$) and the brightest absolute magnitude ($m^{\mathrm{min}}$=-27). These numbers are chosen such that the number of haloes or galaxies above those limits is so small that it does not impact the matching results from SHAM. We create a linearly-spaced vector of negative magnitudes $\bm{\nu}\!\in\!\mathbb{R}^{n}\!:\!\nu_{i}\!=\!-m_{i}$ and logarithmically-spaced halo mass vector $\bm{\mu}\!\in\!\mathbb{R}^{n}\!:\!\mu_{i}\!=\!\log_{10}M_{i}$, both with $n\!=\!500$ within the limits described above. We generate the discretized halo mass function $\mathbf{p}\!\in\!\mathbb{R}_{+}^{n}$ using equations 9 and 11, and the galaxy absolute magnitude function $\mathbf{q}\in\mathbb{R}_{+}^{n}$ using Eqn. 12. We show $\mathbf{p}$ and $\mathbf{q}$ in side panels in Fig. 1. Since the choice of cost metric does not impact the unregularized optimal transport plan, we use $C_{ij}\!=\!(\nu_{i}\!-\!\mu_{j})^{2}$ and obtain the optimal transport plan by solving the problem in Eqn. 4 using linear programming. In the standard SHAM approach, we sample catalogues in an interval $z\!\in\![0.495,0.505]$, which is typical for practical abundance matching, resulting in $\sim\!10^{8}$ objects. The galaxy-halo connection can only be computed up to a mass limit of about $10^{14}M_{\odot}$, as the rapid decline of the mass functions prevents any higher-mass haloes in the simulated catalogue.
@@
The resulting optimal transport plan $\mathbf{Q}$ is presented in Fig. 1 and compared with the standard SHAM approach: they align almost perfectly. Residual differences arise because standard SHAM uses finite simulation catalogues, where massive halos are few and subject to sampling noise – especially at the high-mass end. In contrast, SHAM-OT works with binned distributions, allowing smooth extension to higher halo masses without discreteness effects. Furthermore, we find a computational speed improvement of $O(100)$ for SHAM-OT compared to standard SHAM. SHAM-OT’s computational efficiency allows us to easily study the impact of changes in the luminosity or halo mass function. In Fig. 2, we show how different values of the cosmological and luminosity function parameters impact the galaxy-halo connection.
@@ 6 Entropic Regularization as a Scatter Model
In the OT framework, the marginals of the transport plan are guaranteed to match the input distributions. The effect of scatter can be modelled naturally by the use of regularization of the transport plan: increasing the regularization parameter $\epsilon$ will result in a smoother plan. For the regularized OT, the choice of cost metric directly affects the structure of the resulting transport plan. Various noise model behaviours, including the dependence on halo mass or galaxy luminosity, can therefore be designed through different choices of the cost function, as long as it satisfies the metric requirements.
We demonstrate the capacity of the entropic regularization to model various scattering properties, including ones that mimic the log-normal approach. We use the same analytical prescription with the same cuts for the halo mass and luminosity function as described in Sec. 5, and make sure that potential edge effects at the lower mass limit have negligible impact on the transport plan. We compute the standard SHAM solution with Gaussian scatter in magnitude, using the method from B10 for three values of scatter $\sigma_{\mathrm{mag}}$. To properly emulate this noise behaviour with SHAM-OT, we adapt the basic cost function from the previous section by two parameters $\mu_{\mathrm{sc}}$ and $\alpha$ giving $C_{ij}\!=\!(\nu_{i}\!-\!\mu_{j})^{2}/(\mu_{j}\!-\!\mu_{\mathrm{sc}})^{\alpha}$. We use $\mu_{\mathrm{sc}}\!=\!9.94$ and $\alpha\!=\!-1.7$ for our fiducial setup. The OT solution is obtained with the Sinkhorn algorithm (Alg. 1). Fig. 4 shows the comparison between the two approaches for different noise levels and corresponding $\epsilon$ for our fiducial case. SHAM-OT matches the log-normal model well, both in terms of the mean of the distribution (left panel) and its shape (right panel) and is able to reproduce the behaviour of the standard SHAM.
Beyond emulating the log-normal scatter, regularized OT offers a flexible framework to model different scatter behaviours. Fig. 4 illustrates this flexibility by comparing the conditional transport plan for two different choices of cost functions. Further exploration of cost functions, along with the development of physically motivated choices to reproduce existing scatter behaviour with high precision, is left for future work.
@@ 7 SHAM-OT with haloes from simulations
In practice, the halo mass function is often not analytically given, but can instead be computed from empirical halo distributions based on simulations. We use the approximate code PINOCCHIO (Monaco et al., 2002a; Taffoni et al., 2002; Monaco et al., 2002b, 2013; Munari et al., 2017; Rizzo et al., 2017) to simulate the past light cone of dark matter haloes. PINOCCHIO uses Lagrangian perturbation theory as an approximation which makes it faster than full N-body simulations. We assign subhaloes to the halo catalogue based on the merger tree using the method described in Berner et al. (2022) and label the haloes and subhaloes as hosts for red or blue galaxies based on the model from Berner et al. (2024). This model is based on two key parameters: the halo mass limit $M_{\text{limit}}\!=\!8\!\times\!10^{12}\,h^{-1}M_{\odot}$ and the satellite quenching time $t_{\text{quench}}\!=\!2.0$ Gyr. Central haloes with masses above $M_{\text{limit}}$ host red galaxies, while lower-mass haloes host blue galaxies. Galaxies in subhaloes remain blue until they have spent more than $t_{\text{quench}}$ within their host, after which they transition to red. As a result, we obtain catalogues of haloes and subhaloes with their corresponding mass, position in 3D space, and population label (i.e. red or blue). Note that the methods above can easily be replaced by any other approach that provides a halo catalogue containing the same parameters. We split the host catalogue into redshift bins with width $\Delta z\!=\!0.01$. For each bin, we obtain a host mass function as a histogram, including the upper (heaviest host in the redshift bin) and lower (mass resolution of the simulation) mass limits. We then utilize the luminosity function as defined in F25 with the parameters for the red and blue population from the first index of the posterior distribution. Similar to the procedure described in Sec. 5, we determine the faintest absolute magnitude required such that the number of hosts equals the expected number of galaxies in this redshift bin. Using the constructed distributions for the host mass and galaxy absolute magnitudes, we compute the OT solution for each redshift bin, for both red and blue galaxies. The resulting galaxy-halo relation at different redshifts, for both red and blue galaxies, is shown in Fig. 5. Introducing scatter to this relation would be straightforward, however a careful choice of the cost matrix would then be necessary, see also Section 6.